OpenStudy (anonymous):

can someone solve x^2-6x=7

OpenStudy (saifoo.khan):

x^2 -6x - 7 = 0

OpenStudy (saifoo.khan):

olve!

OpenStudy (saifoo.khan):

Sovle*

OpenStudy (katrinakaif):

Once you get to saif's equation: then think..what mulitplies into -7 and adds into -6 -7 x 1 = -7 -7 + 1 = -6 NOw factor: x^2 -6x - 7 = 0 x^2 -7x)(1x - 7) = 0 x(x - 7) 1( x - 7) = (x + 1)(x - 7) = 0 x + 1 = 0 and x - 7 = 0 x = -1 or x = +7

OpenStudy (anonymous):

x=-1,7